A train has a max speed of 20m/s w/c it can achieve at an acceleration of 0.25m/s². With its brake fully applied, the train has a deceleration of 0.5m/s². What’s the shortest time that the train can travel b/w stations 8k m apart if it stops at both?

1
Expert's answer
2022-04-25T13:33:08-0400

"t=\frac sv=\frac s{v_0-\sqrt{(a_1+a_2)s}+}=7~min."